# Policy Learning with Confidence

### v1

@@ Abstract.
This paper proposes a framework for selecting policies that maximize expected benefit in the presence of estimation uncertainty, by controlling for estimation risk and incorporating risk aversion. The proposed method explicitly balances the size of the estimated benefit against the uncertainty inherent in its estimation, ensuring that chosen policies meet a reporting guarantee, namely that the actual benefit of the implemented policy is guaranteed not to fall below the reported estimate with a pre-specified confidence level. This approach applies to a variety of settings, including the selection of policy rules that allocate individuals to treatments based on observed characteristics, using both experimental and non-experimental data; and the allocation of limited budgets among competing social programs; as well as many others. Across these applications, the framework offers a principled and robust method for making data-driven policy choices under uncertainty. In broader terms, it focuses on policies that are on the efficient decision frontier, describing policies that offer maximum estimated benefit for a given acceptable level of estimation risk.
We thank Isaiah Andrews, Nathan Hendren, Matt Masten, Vira Semenova, and participants at the July 2024 Munich Econometrics Workshop, a Duke microeconometrics breakfast seminar, and a CUNY Graduate Center seminar for helpful comments and discussion. Tian Xie provided excellent research assistance.
@@
Department of Economics, Duke University.
@@ 1. Introduction
Consider a decision-maker (DM) faced with choosing from a menu of policies to maximize expected benefit. The DM could for example be a planner deciding how to allocate treatments to heterogeneous individuals, a firm choosing different potential innovations in which to invest resources, an auctioneer choosing an optimal auction design or reserve price, or a policy-maker choosing expenditure shares to allocate to government programs to maximize a measure of benefit to the public. If the DM knew the benefit that would be obtained from implementing each policy, the choice would be easy; she would then simply choose the policy that would yield the highest benefit. However, the DM lacks such precise knowledge, but instead has data that can be used to consistently estimate the benefit that would be achieved by each policy. The estimates are measured with varying degrees of precision, as reflected by their standard errors. How should the DM use the available information for policy selection?
@@
An intuitively appealing choice is to select the policy with the highest estimated benefit, the so-called plug-in rule, which simply replaces the population objective with its sample analog in determining policy choice. In the literature on statistical treatment rules, this is referred to as empirical welfare maximization (EWM). [^1] See for example Kitagawa and Tetenov (2018) and Athey and Wager (2021). Andrews et al. (2024) alternatively refers to this as the “natural rule” or “picking the winners”. The same logical task applies more broadly to a variety of economic contexts, for example when researchers use structural models, they obtain estimates of the performance of different policies that can then be used to inform policy choice. Manski (2021) calls this “as-if” optimization, which he describes as, “specification of a model, point estimation of its parameters, and use of the point estimate to make a decision that would be optimal if the estimate were accurate.” Here we refer to such rules interchangeably as plug-in rules or empirical welfare maximization (EWM) rules.
@@
Unfortunately, the available estimates for the welfare of different choices are generally not identical to their population values. Estimation error could result in the EWM rule delivering suboptimal policy choice. It may be observed, for example, that among a menu of options policy A has the highest estimated benefit with policy B coming in a close second, but that the standard error of policy B is much smaller than that of policy A, reflecting that the benefit of policy B is estimated much more precisely. Could it be that policy B is actually better than policy A, and that the higher estimated benefit of policy A is simply down to a lack of precision in the estimates? Should the DM consider choosing policy B since it is estimated more precisely to hedge against estimation error?
@@
This paper proposes a framework for selecting policies with the goal of maximizing expected benefit in the presence of estimation uncertainty, explicitly accounting for estimation risk in the decision rule. As inputs to her decision the DM has consistent estimates for the benefit of each available policy, and their corresponding sample variance matrix, with sample estimates approximately normally distributed around their population means. [^2] If the estimates are from independent samples the sample variance matrix will simply be a diagonal matrix with each policy’s sample variance along the diagonal. We define and analyze a class of risk-aware policy rules that provide a principled manner for explicitly balancing the estimated benefit of each policy against its estimation risk. Policies that perform well in sample but are estimated imprecisely may be passed over in favor of less risky policies that are not optimal according to the EWM rule but that can be more confidently assured to achieve high performance. Risk-aware rules select a policy from the (Pareto) efficient decision frontier that balances performance and precision as measured by trading off higher sample estimates against smaller sample variance, where the exact tradeoff is governed by the DM’s choice of policy rule from the class of risk-aware rules. Consequently, such rules select the best performing policy among those within the DM’s choice of acceptable estimation error. They can alternatively be interpreted as selection rules obtained by a risk-averse Bayesian DM who minimizes an expected loss function that takes a weighted average of regret and estimation error. [^3] This is shown in Section 2.1. Such dual-objective loss functions have been explored in the statistics literature in e.g. Gupta and Miescke (1986), Gupta and Miescke (1990), and see also Liese and Miescke (2008), as a way to unify the problems of selection and estimation when sampling from independent normal populations. We establish favorable regret properties for this class of rules, as well as conditions under which they deliver a reporting guarantee, ensuring with high confidence that the actual benefit exceeds a lower threshold.
@@
To illustrate, consider a setting in which a budget-constrained DM must choose how to allocate funds across several social programs that affect the same group of individuals. Suppose the DM wishes to use the Marginal Value of Public Funds (MVPF), spear-headed by Hendren and Sprung-Keyser (2020), for this purpose. The MVPF of each social program is the ratio of its marginal benefit to its net marginal cost to the government, inclusive of the impact of any behavioral responses on the government budget. For instance, an MVPF of 1.5 indicates that each \$1 of net government spending generates \$1.50 in benefits for beneficiaries. For a utilitarian DM, the MVPF provides a metric to compare the “bang for the buck” of different policies that target the same beneficiaries. A DM interested in allocating additional marginal expenditure will prioritize programs with the highest MVPF to maximize welfare. [^4] Here we abstract from distributional incidence and assume the DM values dollars equally across the beneficiaries of each policy. If welfare weights vary across beneficiaries, as in the original framework of Hendren and Sprung-Keyser (2020), our analysis can accommodate this by adjusting each MVPF in the value function using the DM’s designated weights. Example welfare weights have been derived in Hendren (2020).
@@
We propose Policy Learning with Confidence, or PoLeCe, among the class the of risk-aware policy rules that lie on the efficient decision frontier. Like all risk-aware rules, PoLeCe uses observable measures of statistical variation to adjust policy choice relative to a simple plug-in rule by penalizing the observed performance of each policy by an amount proportional to its standard error. Specifically PoLeCe is a risk-aware decision rule that uses a data-dependent penalization factor so that the resulting objective value automatically provides the lower bound of a one-sided confidence interval for the welfare achieved by the selected rule. Thus, the rule delivers a reporting guarantee, ensuring with high confidence that the actual benefit exceeds a lower threshold; no adjustment for post-selection inference is required. Importantly, policies that allocate fractional shares to different treatments or programs can be allowed, producing a richer frontier than that available from singleton allocations such as the individual programs indicated by hollow squares in Figure 1.
@@
In the above example, all that was required to map out the efficient decision frontier, and indeed all that is needed to determine the optimal allocation following our proposal, are point estimates for the benefit of each available policy and their joint variance-covariance. [^6] Our proposed PoLeCe rule uses the sample correlation of the estimates to calibrate the precision penalty to achieve the aforementioned reporting guarantee. The general class of risk-aware decision rules described in Section 2.1 allows for different choices for the penalization factor, which corresponds to the green slope highlighted in Figure 1. Other risk-aware rules may thus only require standard errors rather than estimates of the entire correlation structure, depending on their choice of penalization factor. This can be obtained in applications that feature a variety of different models and sampling processes. An important special case to which our framework applies is the analysis of optimal treatment assignment using data on individual-specific treatments and allocations, such as from an RCT.
@@
The analysis of statistical treatment rules in such settings has received renewed attention in the recent literature, starting with the pioneering work of Manski (2004), who considers treatment choice on the basis of minimax (expected) regret. Follow-up works include a decision theoretic framework introduced by Dehejia (2005), finite-sample bounds considered by Stoye (2009), and an asymptotic framework introduced by Hirano and Porter (2009, 2020). Kitagawa and Tetenov (2018) proposes empirical welfare maximization (EWM) rules for experimental data, which selects a policy with the highest estimated welfare (empirical welfare), thus constituting a plug-in rule for treatment effect models with experimental data. Kitagawa and Tetenov (2018) proves the optimality of EWM rules in the sense that as the sample size increases expected regret converges to zero at the minimax rate. Athey and Wager (2021) and Demirer et al. (2019) propose doubly-robust estimation of average welfare, in which case this result obtains even with quasi-experimental data. In the context of this literature, PoLeCe offers a modification of EWM that incorporates estimation risk into the decision rule. The EWM rule corresponds to the PoLeCe rule with penalization factor of zero on the standard error, putting no weight on estimation risk. Inference approaches for optimal treatment assignment and/or welfare complementary to EWM, but which do not propose alternative assignment rules to account for uncertainty, include those of Andrews et al. (2024), Armstrong and Shen (2023), Ponomarev and Semenova (2024), and Rai (2019), as well as the online supplement of the original paper by Kitagawa and Tetenov (2018).
@@
The literature on statistical treatment rules features some recent alternative proposals to account for estimation uncertainty in the decision rule. In settings with a budget constraint, Sun (2021) proposes a rule that penalizes policies with large uncertainty with meeting the constraint. Kitagawa et al. (2022) focuses on the variability of regret and shows that the EWM rule is dominated by fractional rules when the objective function is the sum of squared expected regret and the variance of regret. If the estimated welfare is independent across policies, then one may appeal to the literature on optimal selection such as Gupta and Miescke (1986), referenced above. However, in the context of policy learning, since the groups assigned by different policies might not be mutually exclusive, the estimated welfare associated with different policies may not be independent. For instance, when treatment allocation is determined by an income threshold, the welfare estimates for groups falling below different thresholds will be correlated by construction.
@@
The paper proceeds as follows. Section 2 sets out our approach to risk-aware policy choice that selects policies from the estimated efficient decision frontier. Section 2.1 motivates the approach using selection of policies for treatment choice and investment allocation in public programs as key illustrative examples. Section 2.2 considers the general problem faced by a decision-maker concerned with estimation risk, leading to risk-aware policy rules. Section 2.3 develops favorable regret and lower confidence guarantees for the general class of risk-aware policy rules, and Section 2.4 formally sets out the proposed PoLeCe rule as the risk-aware rule that maximizes a lower confidence bound on maximal welfare. Section 2.5 discusses refinements and further theoretical points and Section 2.6 provides evidence from Monte Carlo experiments calibrated to an empirical application on the use of different methods for encouraging vaccination uptake. Section 3 provides four empirical applications that demonstrate the application of PoLeCe, and Section 4 concludes.
@@ 2.1. Motivation and Risk-Aware Policy Choice
The value function of a policy is then defined as
@@
Under standard conditions, including $\mathrm{E}[H^{2}(\pi)]<\infty$ – a relevant overlap condition [^7] This requires that policies do not place weight on treatments for which there is insufficient randomization at particular values of $X$. Note that the standard overlap condition can fail, but the value of a particular policy is identified, for example, if $H(\pi)$ and $Y$ have a finite second moment. – we obtain the following regression and representer formulas for identifying the policy value:
@@
which can be viewed either as aggregating up regressions or as reweighting observed outcomes by how the policy $\pi$ differs from the propensity score $p$. Combining these two approaches leads to efficient representation and estimation of the value function via the influence function (e.g., Newey, 1994):
@@
$\psi_{\pi}(W):=\sum_{t\in\mathcal{T}}(m(t,X)-m(0,X))\pi(t\mid X)+H(\pi)\,\bigl{[}Y-m(T,X)\bigr{]}$ (3)
where $m(t,X)\coloneqq\mathrm{E}\left[Y\mid T=t,X\right]$ and $H(\pi)=\pi(t\mid X)\ \left[\frac{\mathbf{1}\{T=t\}}{p(t\mid X)}-\frac{\mathbf{1}\{T=0\}}{p(0\mid X)}\right]$.
We follow the convention of encapsulating in $X$ both convariates that are confounders as well as those for which the planner may assign different treatments.
@@
Different treatment policies $\pi$ can have markedly different reward and risk profiles, and some may be much riskier than others. For example, policies that “turn on” treatments with near-zero propensities may be especially risky, since
@@
The decision maker has to choose an optimal policy in $\Pi$ based on data $\widehat{V}(\pi)$ that is (approximately) Gaussian. We argue that an optimal risk-aware decision can be achieved by maximizing the empirical policy value offset by the estimated risk times a critical value:
$\widehat{\pi}_{RW}\;\in\;\arg\max\Bigl{\{}\widehat{V}(\pi)-k\,\widehat{s}(\pi):\pi\in\Pi\Bigr{\}},$ (4)
where $k$ is a critical value (depending on risk preferences) and $\widehat{s}(\pi)$ is a consistent estimator of the estimation risk $s(\pi)$. By varying $k$, we trace out the efficient decision frontier of treatment policies, as illustrated by Figure 1 in the introduction. Each point on the frontier corresponds to a particular $k>0$. We provide fast algorithms in Appendix B to efficiently compute the frontier for any application.
Our leading proposal to choose $k$ is to meet certain reporting guarantees with high confidence, thereby conducting “policy learning with confidence” (PoLeCe). [^9] Other choices of $k$ are of course possible and can be rationalized as optimal under risk aversion. We discuss those later in the paper. The resulting choice generally differs from the empirical welfare maximizer (EWM):
$\widehat{\pi}_{EWM}\;\in\;\arg\max\{\widehat{V}(\pi):\pi\in\Pi\}.$ (5)
The EWM rule arises when risk aversion vanishes, i.e. when $k\searrow 0$. We provide a comparative analysis of these two rules in what follows. Before we present details of our PoLeCe proposal, we discuss another class of applications.
@@
which is the unit simplex intersected with a rectangular set of constraints. These additional constraints may reflect diversity or other requirements that allocations must satisfy. The value of the allocation $\pi$ is
@@
so that the estimated values of the allocations are also approximately Gaussian:

### v2

@@ Abstract.
This paper introduces a framework for selecting policies that maximize expected welfare under estimation uncertainty. The proposed method explicitly balances the size of the estimated welfare against the uncertainty inherent in its estimation, ensuring that chosen policies meet a reporting guarantee, namely, that actual welfare is guaranteed not to fall below the reported estimate with a pre-specified confidence level. We produce the efficient decision frontier, describing policies that offer maximum estimated welfare for a given acceptable level of estimation risk. We apply this approach to a variety of settings, including the selection of policy rules that allocate individuals to treatments and the allocation of limited budgets among competing social programs.
We thank Isaiah Andrews, Nathan Hendren, Matt Masten, Whitney Newey, Vira Semenova, Jesse Shapiro, Aleksey Tetenov and conference and seminar participants at the Munich Econometrics Workshop, Duke University, CUNY, University of Geneva, Tilburg University, the Encounters in Econometric Theory workshop, CUHK Shenzhen, and IAAE for helpful comments and discussion. Tian Xie provided excellent research assistance. Liyang Sun acknowledges generous support from the Economic and Social Research Council (new investigator grant UKRI607).
@@
§Department of Economics, Duke University.
@@
Keywords: budget allocation, risk-aware policy learning, statistical decision theory
JEL classification codes: C14, C44, C52.
@@ 1. Introduction
Consider a decision-maker (DM) faced with choosing from a menu of policies to maximize expected welfare. The DM could for example be a planner deciding how to allocate treatments to heterogeneous individuals, a firm choosing different potential innovations in which to invest resources, an auctioneer choosing an optimal auction design or reserve price, or a policy-maker choosing expenditure shares to allocate to government programs to maximize a measure of welfare to the public. If the DM knew the welfare that would be obtained from implementing each policy, the choice would be easy; she would then simply choose the policy that would yield the highest welfare. However, the DM lacks such precise knowledge, but instead has data that can be used to consistently estimate the welfare that would be achieved by each policy. The estimates are measured with varying degrees of precision, as reflected by their standard errors. How should the DM use the available information for policy selection?
@@
An intuitively appealing choice is to select the policy with the highest estimated welfare, the so-called plug-in rule, which simply replaces the population objective with its sample analog in determining policy choice. In the literature on statistical treatment rules, this is referred to as empirical welfare maximization (EWM). [^1] See for example Kitagawa and Tetenov (2018) and Athey and Wager (2021). Andrews et al. (2024) alternatively refers to this as the “natural rule” or “picking the winners”. The same logical task applies more broadly to a variety of economic contexts, for example when researchers use structural models, they obtain estimates of the performance of different policies that can then be used to inform policy choice. Manski (2021) calls this “as-if” optimization, which he describes as, “specification of a model, point estimation of its parameters, and use of the point estimate to make a decision that would be optimal if the estimate were accurate.” Here we refer to such rules interchangeably as plug-in rules or empirical welfare maximization (EWM) rules.
@@
Unfortunately, the available estimates for the welfare of different choices are generally not identical to their population values. Estimation error could result in the EWM rule delivering suboptimal policy choice. It may be observed, for example, that among a menu of options policy A has the highest estimated welfare with policy B coming in a close second, but that the standard error of policy B is much smaller than that of policy A, reflecting that the welfare of policy B is estimated much more precisely. Could it be that policy B is actually better than policy A, and that the higher estimated welfare of policy A is simply down to a lack of precision in the estimates? Should the DM consider choosing policy B since it is estimated more precisely to hedge against estimation error?
@@
This paper proposes a framework for selecting policies with the goal of maximizing expected welfare in the presence of estimation uncertainty, explicitly accounting for estimation risk in the decision rule. As inputs to her decision the DM has consistent estimates for the welfare of each available policy, and their corresponding sample variance matrix, with sample estimates approximately normally distributed around their population means. [^2] If the estimates are from independent samples the sample variance matrix will simply be a diagonal matrix with each policy’s sample variance along the diagonal. We define and analyze a class of risk-aware policy rules that provide a principled manner for explicitly balancing the estimated welfare of each policy against its estimation risk. Policies that perform well in sample but are estimated imprecisely may be passed over in favor of less risky policies that are not optimal according to the EWM rule but that can be more confidently assured to achieve high performance. Risk-aware rules select a policy from the (Pareto) efficient decision frontier that balances performance and precision as measured by trading off higher sample estimates against smaller sample variance, where the exact tradeoff is governed by the DM’s choice of policy rule from the class of risk-aware rules. Consequently, such rules select the best performing policy among those within the DM’s choice of acceptable estimation error. They can alternatively be interpreted as selection rules obtained by a risk-averse Bayesian DM who minimizes an expected loss function that takes a weighted average of regret and estimation error. [^3] This is shown in Section 2.5. Such dual-objective loss functions have been explored in the statistics literature in e.g. Gupta and Miescke (1986), Gupta and Miescke (1990), and see also Liese and Miescke (2008), as a way to unify the problems of selection and estimation when sampling from independent normal populations. We establish favorable regret properties for this class of rules, as well as conditions under which they deliver a reporting guarantee, ensuring with high confidence that the actual welfare exceeds a lower threshold.
@@
To illustrate, consider a setting in which a budget-conscious DM needs to allocate funds across several social programs that affect the same group of individuals, either in response to a slight budget increase or for incremental cost-cutting. Suppose the DM wishes to use the Marginal Value of Public Funds (MVPF), spear-headed by Hendren and Sprung-Keyser (2020), for this purpose. The MVPF of each social program is the ratio of its marginal benefit to its net marginal cost to the government, inclusive of the impact of any behavioral responses on the government budget. For instance, an MVPF of 1.5 indicates that each \$1 of net government spending generates \$1.50 in benefits for beneficiaries. For a utilitarian DM, the MVPF provides a metric to compare the “bang for the buck” of different policies that target the same beneficiaries. A DM interested in allocating additional marginal expenditure will prioritize programs with the highest MVPF to maximize welfare. [^4] Here we abstract from distributional incidence and assume the DM values dollars equally across the beneficiaries of each policy. If welfare weights vary across beneficiaries, as in the original framework of Hendren and Sprung-Keyser (2020), our analysis can accommodate this by adjusting each MVPF in the value function using the DM’s designated weights. Example welfare weights have been derived in Hendren (2020). Other metrics for measuring program’s marginal benefit and cost have also been discussed in García and Heckman (2022).
@@
We propose Policy Learning with Confidence, or PoLeCe, among the class of risk-aware policy rules that lie on the efficient decision frontier. Like all risk-aware rules, PoLeCe uses observable measures of statistical variation to adjust policy choice relative to a simple plug-in rule by penalizing the observed performance of each policy by an amount proportional to its standard error. Specifically PoLeCe is a risk-aware decision rule that uses a data-dependent penalization factor so that the resulting objective value automatically provides the lower bound of a one-sided confidence interval for the welfare achieved by the selected rule. Thus, the rule delivers a reporting guarantee, ensuring with high confidence that the actual welfare exceeds a lower threshold; no adjustment for post-selection inference is required. Importantly, policies that allocate fractional shares to different treatments or programs can be allowed, producing a richer frontier than that available from singleton allocations such as the individual programs indicated by hollow squares in Figure 1.
@@
In the above example, all that was required to map out the efficient decision frontier, and indeed all that is needed to determine the optimal allocation following our proposal, are point estimates for the welfare of each available policy and their joint variance-covariance. [^6] Our proposed PoLeCe rule uses the sample correlation of the estimates to calibrate the precision penalty to achieve the aforementioned reporting guarantee. The general class of risk-aware decision rules described in Section 2.1 allows for different choices for the penalization factor, which corresponds to the green slope highlighted in Figure 1. Other risk-aware rules may thus only require standard errors rather than estimates of the entire correlation structure, depending on their choice of penalization factor. This can be obtained in applications that feature a variety of different models and sampling processes. An important special case to which our framework applies is the analysis of optimal treatment assignment using data on individual-specific treatments and allocations, such as from an RCT. For example, in a real-world setting of treatment assignment, while GiveWell uses EWM based on RCT results, they have recently highlighted the importance of accounting for estimation uncertainty in decision-making for the sake of “transparency” (Salisbury, 2024). Our proposal offers a practical solution to this concern.
The analysis of statistical treatment rules in such settings has received renewed attention in the recent literature, starting with the pioneering work of Manski (2004), who considers treatment choice on the basis of minimax (expected) regret. Follow-up works include a decision theoretic framework introduced by Dehejia (2005), finite-sample bounds considered by Stoye (2009), and an asymptotic framework introduced by Hirano and Porter (2009, 2020). Hypothesis testing is a conventional approach to decision-making under uncertainty. As argued by Tetenov (2012), it can be interpreted as a DM having asymmetric aversion to Type I and Type II regret.
@@
Kitagawa and Tetenov (2018) proposes empirical welfare maximization (EWM) rules for experimental data, which is an “as-if” optimization rule that selects a policy with the highest estimated welfare (empirical welfare), thus constituting a plug-in rule for treatment effect models with experimental data. Kitagawa and Tetenov (2018) proves the optimality of EWM rules in the sense that as the sample size increases expected regret converges to zero at the minimax rate. Athey and Wager (2021) and Demirer et al. (2019) propose doubly-robust estimation of average welfare, in which case this result obtains even with quasi-experimental data. In the context of this literature, PoLeCe offers a modification of EWM that incorporates estimation risk into the decision rule. The EWM rule corresponds to the PoLeCe rule with penalization factor of zero on the standard error, putting no weight on estimation risk. Inference approaches for optimal treatment assignment and/or welfare complementary to EWM, but which do not propose alternative assignment rules to account for uncertainty, include those of Armstrong and Shen (2023), Ponomarev and Semenova (2024), and Rai (2019), as well as the online supplement of the original paper by Kitagawa and Tetenov (2018). As an extension of EWM, Mbakop and Tabord-Meehan (2021) proposes a penalized welfare maximization rule that penalizes for the complexity of the policy, and establishes an oracle property for model selection.
Andrews and Chen (2025) proposes the use of certified decisions, which comprise the pairing of a decision rule with an accompanying inferential guarantee. Their results provide a decision theoretic justification for the PoLeCe rule, showing that it belongs to an essentially complete class of decision rules for a DM who requires a coverage guarantee, and therefore dominates the pairing of EWM with the studentized projection approach in Andrews et al. (2024).
@@
PoLeCe also contributes to a recent literature featuring some alternative proposals to directly account for estimation uncertainty in the decision rule. In settings with a budget constraint, Sun (2021) proposes a rule that penalizes policies with large uncertainty in meeting the constraint, and Moon (2025) proposes using empirical Bayes estimates, leveraging shared information across related policies. Kitagawa et al. (2022) focuses on the variability of regret and shows that the EWM rule is dominated by fractional rules when the objective function is the sum of squared expected regret and the variance of regret. If the estimated welfare is independent across policies, then one may appeal to the literature on optimal selection such as Gupta and Miescke (1986), referenced above. However, in the context of policy learning, since the groups assigned by different policies might not be mutually exclusive, the estimated welfare associated with different policies may not be independent. For instance, when treatment allocation is determined by an income threshold, the welfare estimates for groups falling below different thresholds will be correlated by construction.
@@
The paper proceeds as follows. Section 2 sets out our approach to risk-aware policy choice that selects policies from the estimated efficient decision frontier. Section 2.1 motivates the approach using selection of policies for treatment choice and investment allocation in public programs as key illustrative examples. Section 2.2 develops favorable regret and lower confidence guarantees for the general class of risk-aware policy rules, and Section 2.3 formally sets out the proposed PoLeCe rule as the risk-aware rule that maximizes a lower confidence bound on maximal welfare. Section 2.4 provides evidence from Monte Carlo experiments calibrated to empirical applications. Section 2.5 provides alternative motivations for risk-aware decisions. Section 2.6 discusses refinements and further theoretical points. Section 3 provides four empirical applications that demonstrate the application of PoLeCe, and Section 4 concludes.
@@ 2.1. Motivation and Risk-Aware Policy Choice
The welfare function of a policy is then defined as
@@
Under standard conditions, including $\mathrm{E}[H^{2}(\pi)]<\infty$ – a relevant overlap condition [^7] This requires that policies do not place weight on treatments for which there is insufficient randomization at particular values of $X$. Note that the standard overlap condition can fail, but the welfare of a particular policy is identified, for example, if $H(\pi)$ and $Y$ have a finite second moment. – we obtain the following regression and representer formulas for identifying the policy’s welfare:
@@
which can be viewed either as aggregating up regressions or as reweighting observed outcomes by how the policy $\pi$ differs from the propensity score $p$. Combining these two approaches leads to efficient representation and estimation of the welfare function via the influence function (e.g., Newey, 1994):
@@
$\psi_{\pi}(W):=\sum_{t\in\mathcal{T}}(g(t,X)-g(0,X))\pi(t\mid X)+H(\pi)\,\bigl{[}Y-g(T,X)\bigr{]}$ (3)
where we re-define $H(\pi)=\sum_{t\in\mathcal{T}}\ \left[\frac{\mathbf{1}\{T=t\}}{p(t\mid X)}-\frac{\mathbf{1}\{T=0\}}{p(0\mid X)}\right]\pi(t\mid X)$.
We follow the convention of encapsulating in $X$ both covariates that are confounders as well as those for which the planner may assign different treatments.
@@
Different treatment policies $\pi$ can have markedly different welfare and estimation risk profiles, and some may be much riskier than others. For example, policies that “turn on” treatments with near-zero propensities may be especially risky, since
@@
The decision maker has to choose an optimal policy in $\Pi$ based on data $\widehat{V}(\pi)$ that is (approximately) Gaussian. We argue that an optimal risk-aware decision can be achieved by maximizing the empirical welfare offset by the estimation risk times a critical value:
$\widehat{\pi}_{\text{RW}}\;\in\;\arg\max\Bigl{\{}\widehat{V}(\pi)-k\,\widehat{s}(\pi):\pi\in\Pi\Bigr{\}},$ (4)
where $k$ is a critical value and $\widehat{s}(\pi)$ is a consistent estimator of the estimation risk $s(\pi)$. By varying $k$, we trace out the efficient decision frontier of treatment policies, as illustrated by Figure 1 in the introduction. Each point on the frontier corresponds to a particular $k>0$. We provide fast algorithms in Appendix B to efficiently compute the frontier for any application.
Our leading proposal to choose $k$ is to meet certain reporting guarantees with high confidence, thereby conducting “policy learning with confidence” (PoLeCe). [^9] Other choices of $k$ are of course possible and can be rationalized as optimal under risk aversion. We discuss those later in Section 2.5. The resulting choice generally differs from the empirical welfare maximizer (EWM):
$\widehat{\pi}_{\text{EWM}}\;\in\;\arg\max\{\widehat{V}(\pi):\pi\in\Pi\}.$ (5)
The EWM rule arises when $k\searrow 0$. We provide a comparative analysis of these two rules in what follows. Before we present details of our PoLeCe proposal, we discuss another class of applications.
@@
which is the unit simplex intersected with a rectangular set of constraints. These additional constraints may reflect diversity or other requirements that allocations must satisfy. The welfare of the allocation $\pi$ is
@@
so that the estimated welfare of the allocations are also approximately Gaussian: